OpenStudy (anonymous):

Given the function g(x) = 4x − 5, compare and contrast g(2) and g(−4). Choose the statement that is true concerning these two values.

hero (hero):

g(2) = 4(2)-5 = 8 - 5 = 3 g(-4)=(-4)-5 = -4 + -5 = -9

OpenStudy (anonymous):

thanks u